OpenStudy (anonymous):

what is 4w in simplified from

OpenStudy (baldymcgee6):

That is simplified form.

OpenStudy (anonymous):

\[\sqrt{4w}what is \it simplifed\]

OpenStudy (baldymcgee6):

Well you can write this as\[\sqrt{4w}=2\sqrt{w}\]

OpenStudy (anonymous):

\[what is the \square \root of 64\]

OpenStudy (baldymcgee6):

8

OpenStudy (anonymous):

\[\sqrt{45} what i this simplifed\]

OpenStudy (baldymcgee6):

post this in a new question, and it is common courtesy to give out best responses on this site.